• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Jixie Zhang

    Log entry time 00:07:55 on February11,2011

    Entry number 346336

    keyword=Swing Shift Summary

    16:00 Shift started. Target operation trainning was going on.
    17:30 Target operation training finished. Jian-ping checked the Loop1
    target temperature sensors and reset the alarm threshold for both loops. He found some problem about these sensors in Loop1. Check log entry 346309 for details.
    20:30 Hall was closed for sweeping
    20:45 Ramped up bigbite magnet to 518 A. Ramped up HRS magnets, set left arm momentum to 2034 and righ arm to 2190.
    21:25 MCC called saying that the radcon can not sweep the spot behine
    the left arm Q1 power supply box. There was a "High Voltage, Danger!" sign posted there which block the way. We need to take care of the post and redo a sweep by then.
    21:50 Retuened to Restricted access
    22:10 Alex cleared the sign and close the hall for sweeping
    23:30 Finished sweeping the hall.
    23:45 Asked MCC for beam.